OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 1) and (9, 7)?

OpenStudy (anonymous):

\(\dfrac{y_2-y_1}{x_2-x_1}\) thats the slope formula

OpenStudy (anonymous):

With your given numbers (so (1,1) and (9,7)) the first two numbers will be x1 and y1. The next two numbers will be x2 y2. Use the slope formula y2-y1/x2-x1 to determine the answer.

OpenStudy (anonymous):

so you would get 7-1/9-1= 6/8. Simplified it would be 3/4.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

\(\dfrac{7-1}{9-1}\)